OpenStudy (hkmiu):

inverse of arctan (x + pi/2)?

OpenStudy (ksaimouli):

\[\tan^{-1} (x+\frac{ \pi }{ 2 })\]

OpenStudy (jdoe0001):

\(\bf tan(\theta)\quad inverse\implies tan^{-1}(\theta)\\ \quad \\ tan^{-1}(\theta)\quad inverse\implies tan(\theta)\\ \quad \\ \quad \\ tan^{-1}\left(x+\frac{\pi}{2}\right)\quad inverse\implies \square ?\)

OpenStudy (hkmiu):

so I just switch arctan?

OpenStudy (jdoe0001):

I'd think so, yes, since \(\bf tan^{-1}\) is the inverse of \(\bf tan\) and the other way around

OpenStudy (hkmiu):

okay thank you!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

Careful that's not right lmao.